You have a $9,000 portfolio which is invested in stocks A, B, and a risk-free asset. $4,000 is invested in stock A. Stock A has a beta of 1.84 and stock B has a beta of 0.68. How much needs to be invested in stock B if you want a portfolio beta of.95?

Multiple Choice

  • $0

  • $1,750

  • $5,000

  • $3,279

  • $7,279
Answer #1

Portfolio Beta = Weighted avg beta of securites.

Beta of Risk free assets is 0.

Let "X" be the weight in Stock B

Weight in STock A = 4000 / 9000

= 0.4444

Weight in RIsk free Asset = 1 - 0.4444 -X

= 0.5556 - X

Partciculars Weight Beta Wtd Beta
Stock A 0.4444 1.84 0.817696
Stock B X 0.68 0.68X
Risk free Asset 0.5556-X 0 0
Portfolio Beta 0.68X + 0.8177

Thus 0.95 = 0.68X + 0.8177

0.68X = 0.95 - 0.8177

= 0.1323

X = 0.1323 / 0.68

= 0.1946

Investment in B= TotalPortfolio * Weight

= $ 9000 * 0.1946

= $ 1751

OPtion B is correct

Pls comment, if any further assistance is required